Covert 39°C to Farenheit


1
Expert's answer
2021-03-18T16:52:19-0400

We can convert Celsius to Farenheit as follows:


"T_F=T_C\cdot\dfrac{9}{5}+32=39^{\circ}\cdot\dfrac{9}{5}+32=102.2\ ^{\circ}F."